# RASPNet: A Benchmark Dataset for Radar Adaptive Signal Processing Applications

### v2

@@ 3 Dataset Organization
Algorithm 2 : Ordering RASPNet Scenarios by Difficulty
0:  $S$ (set of indices), $b$ (index of baseline scenario)
0:  $Q$ (sequence of indices ordered by $\mathcal{E}(Z^{i},Z^{b})$)
1:  Initialize $F\leftarrow\mathbf{0}_{M}$Initialize the array of $\mathcal{E}$-statistics}
2:  for $i\in S$ do
3:     $F_{i}\leftarrow\mathcal{E}(Z^{i},Z^{b})$
4:  end for
5:  $Q\leftarrow\text{argsort}(F,\leq)$Sort indices in ascending order}
6:  return $Q$
@@
We categorize the $M$ scenarios in RASPNet into four subsets by difficulty: Basic scenarios, for $q_{i}\in\{1,2,\ldots,25\}$, Intermediate scenarios, for $q_{i}\in\{26,27,\ldots,50\}$, Challenging scenarios, for $q_{i}\in\{51,52,\ldots,75\}$, and Eye-watering scenarios, where $q_{i}\in\{76,77,\ldots,100\}$. These subsets are depicted in Figure 3 — the basic scenarios are largely concentrated in the Eastern United States and plains regions, having simpler clutter characteristics (e.g., the Bonneville Salt Flats). The intermediate scenarios have more uneven terrain than the basic scenarios, and encompass hillside regions, coastlines and valleys. The challenging scenarios are largely concentrated in the Western United States, comprising mountainous and variable terrain — this shift to more rugged and uneven terrain introduces clutter characteristics that deviate substantially from the baseline scenario. Lastly, the eye-watering scenarios are composed almost exclusively of mountainous terrain, canyons, and other diversified land types, representing scenarios with highly irregular and unpredictable clutter characteristics.
@@ 4 Dataset Applications
RASPNet has several applications relevant to the radar community. We describe two key applications in this section. Firstly, RASPNet can be used to benchmark data-driven algorithms for target localization against existing approaches. Secondly, RASPNet supports transfer learning in realistic radar scenarios, enabling the adaptation of pre-trained models to new environments — we demonstrate this via a toy example. Across both of these applications, we consider five example scenarios in RASPNet: Bonneville Salt Flats, UT ($q_{i}=1,i=29$) as the baseline scenario, Central Kansas, KS ($q_{i}=4,i=60$) from the Basic scenarios, the Ouachita Mountains, OK ($q_{i}=35,i=62$) from the Intermediate scenarios, the Great Smoky Mountains, TN ($q_{i}=57,i=76$) from the Challenging scenarios, and the Grand Canyon, AZ ($q_{i}=92,i=35$) from the Eye-watering scenarios, wherein $\smash{\tilde{S}=\{29,35,60,62,76\}}$. We first preprocess the RASPNet data via conventional RASP techniques, per Section 4.1. We compiled all results on a NVIDIA GeForce RTX 3090 GPU.
@@ 4.1 Data Preprocessing
To prepare RASPNet for the target localization and transfer learning tasks, we consider sub-regions from the scenarios, $\smash{i\in\tilde{S}}$, which we call radar processing regions. The minimum and maximum range of each radar processing region from the airborne radar platform and the azimuth angle bounds of each radar processing region are provided in Table 3 (see Appendix Section C.1), and the range bin indices are $g\in\{75,76,\ldots,95\}\subset\{1,2,\ldots,G\}$, whereby $G^{\prime}=21$. We run a set of $N^{\prime}=1000$ independent experiments, where in each experiment, a point target is placed in range and azimuth, uniformly at random inside each radar processing region. We select the radar cross section (RCS) of each target independently at random from the uniform distribution, $\mathcal{U}[15,25]$ (dBsm), and we generate $K^{\prime}=100$ independent random realizations of the radar return for each target placement. We consider stationary targets in this example, with $\Lambda^{\prime}=1$, wherein there is no Doppler processing. For compatibility with our generated target data, we divide the $K=10{,}000$ clutter realizations provided by RASPNet for scenario $\smash{i\in\tilde{S}}$ into $N^{*}=100$ separate clutter datasets with $K^{\prime}=100$ realizations, where $K=N^{*}K^{\prime}$. As $\Lambda^{\prime}=1$, we only consider the first pulse from the clutter realizations. For the CVNN target localization benchmark task, we solely consider the first $5$ of the $100$ realizations, wherein $K^{\prime}=5$. Using this construction, we now consider the signal model outlined in Section C.2 of the Appendix.
@@
For both non-CVNN target localization and transfer learning tasks, we leverage datasets of heatmap matrices (Venkatasubramanian et al., 2023), which comprise test statistics obtained via the signal model from Appendix Section C.2, where $\phi=0$. To produce each heatmap matrix, we sweep the radar steering vector, $\mathbf{\tilde{a}}(\theta,0,0)$, across azimuth with step size $\Delta\theta=0.4^{\circ}$, from $\theta_{\text{min}}$ to $\theta_{\text{max}}$, producing an NAMF azimuth spectrum at each range bin. Arranging these spectra across the $G^{\prime}$ range bins gives a size $G^{\prime}\times(\lfloor\frac{\theta_{\text{max}}-\theta_{\text{min}}}{\Delta\theta}\rfloor+1)$ matrix of NAMF test statistics. The values for $\theta_{\text{min}}$ and $\theta_{\text{max}}$ are in Table 3 for scenarios $\smash{i\in\tilde{S}}$. The cumulative dataset contains $N=N^{\prime}N^{*}=100{,}000$ heatmap matrices, formed using all combinations of $\mathbf{X}^{i}[g]$ and $(\mathbf{Z}^{i}[g]+\mathbf{A}^{i}[g])$. For the CVNN target localization task, we consider $3$-dimensional feature vectors, which are formed from the radar return data matrices described in Section C.2 of the Appendix, where $\phi=0$. The cumulative dataset comprises the first $25{,}000$ feature vectors of the $100{,}000$ total combinations.
@@
Figure 4: Test performance comparison on RASPNet scenarios $i\in\tilde{S}$ using CVNN, RVNN, Steinmetz neural network, and analytic neural network. The x-axis represents the training epochs, while the y-axis indicates the target localization performance.
4.2 Target Localization Benchmark
Having obtained the datasets of heatmap matrices for scenarios $\smash{i\in\tilde{S}}$, we benchmark the non-CVNN target localization performance of two existing algorithms — we consider the classical peak-cell midpoint algorithm outlined in (Venkatasubramanian et al., 2022), and compare it with the regression CNN architecture introduced in (Venkatasubramanian et al., 2023). We clarify that other classical algorithms exist for target localization, including gradient-based local search methods; these algorithms can also be benchmarked using RASPNet. The considered regression CNN architecture is further described in Section D.1 of the Appendix.
[…]
$Err(\mathbf{s}_{t}^{*i},\mathbf{\hat{\text{$\mathbf{s}$}}}\vphantom{s}_{t}^{i})=\frac{\sum_{t=1}^{N_{test}}\|\mathbf{s}_{t}^{*i}-\mathbf{\hat{\text{$\mathbf{s}$}}}\vphantom{s}_{t}^{i}\|_{2}}{N_{test}}.$ (5)
[…]
Table 1: Benchmarking target localization performance for $\smash{i\in\tilde{S}}$.
$\boldsymbol{q_{i}}$ $\boldsymbol{i}$ $\boldsymbol{Err_{\text{NAMF}}}$ (m) $\boldsymbol{Err_{\text{CNN}}}$ (m) $\boldsymbol{\mathcal{E}}$-stat.
1 29 21.62 $\pm$ 0 9.88 $\pm$ 0.50 0
4 60 21.83 $\pm$ 0 14.60 $\pm$ 0.82 3.51e-8
35 62 187.31 $\pm$ 0 68.31 $\pm$ 2.23 2.63e-7
57 76 360.53 $\pm$ 0 145.34 $\pm$ 5.91 7.91e-7
92 35 1004.20 $\pm$ 0 220.45 $\pm$ 9.67 4.04e-6
Figure 5: Target localization performance for scenarios $\smash{i\in\tilde{S}\setminus\{29\}}$ after transfer learning $([Err_{\text{CNN}}]_{\text{FT}})$ fine-tuning $([Err_{\text{CNN}}]_{\text{FT}})$.
4.2.1 CVNN Target Localization Benchmark
[…]
@@ 4.3 Transfer Learning
Apart from benchmarking the performance of algorithms, RASPNet can also be leveraged for transfer learning applications in realistic adaptive radar processing scenarios. We recall that the RFView®-generated data exhibits some degree of statistical similarity to real-world data, as examined in Section 2.4. As such, suppose we have collected $K$ clutter realizations, $\{z_{j}^{\omega},\ \forall j\in D\}$ from some real-world scenario, $\omega$, where $z_{j}^{\omega}\sim Z^{\omega}$. Suppose we also have $100$ neural networks trained for some predictive adaptive radar processing task on each of the $100$ RASPNet scenarios, $i\in S$. Our aim is to determine which of the $100$ trained neural networks should be chosen to achieve the best possible performance for the same task on the real-world scenario, $\omega$. We can proceed by ranking the RASPNet scenarios, $i\in S$, by their similarity to scenario $b=\omega$, using Algorithm 2. We hypothesize that the scenario with the lowest $\mathcal{E}$-statistic, relative to the real-world scenario, is the optimal choice for selecting the corresponding trained neural network. To improve the performance of this trained neural network on the real-world scenario, we can use a small number of samples from the real-world scenario to fine-tune it for the task at hand.
[…]

### v3

@@ 3 Dataset Organization
Input: $S$ (set of indices), $b$ (index of baseline scenario)
Output: $Q$ (sequence of indices ordered by $\mathcal{E}(Z^{i},Z^{b})$)
1 $F\leftarrow\mathbf{0}_{M}$ (Initialize the array of $\mathcal{E}$-statistics)
2 for $i\in S$ do $F_{i}\leftarrow\mathcal{E}(Z^{i},Z^{b})$ ;
3 $Q\leftarrow\text{argsort}(F,\leq)$ (Sort the indices of $F$ in ascending order)
return $Q$
Algorithm 2 Ordering RASPNet Scenarios by Difficulty with the $\mathcal{E}$-statistic
@@ 3 $Q\leftarrow\text{argsort}(F,\leq)$ (Sort the indices of $F$ in ascending order)
We categorize RASPNet scenarios into four subsets by difficulty: Basic scenarios: $q_{i}\in\{1,2,\ldots,25\}$, Intermediate scenarios: $q_{i}\in\{26,27,\ldots,50\}$, Challenging scenarios: $q_{i}\in\{51,52,\ldots,75\}$, and Eye-watering scenarios: $q_{i}\in\{76,77,\ldots,100\}$. These four distinct subsets are depicted in Figure 3 — the basic scenarios are largely concentrated within the Eastern United States and the plains regions, having simpler clutter characteristics (e.g., the Bonneville Salt Flats, UT). The intermediate scenarios have more uneven terrain than the basic scenarios, and encompass hillside regions, coastlines and valleys. The challenging scenarios are largely concentrated in the Western United States, comprising mountainous and variable terrain. Lastly, the eye-watering scenarios are composed almost exclusively of mountainous terrain, canyons, and other diversified land types.
@@ 4 Dataset Applications
Among the various applications of RASPNet to the radar and CVNN communities, we describe three key applications in this section: benchmarking data-driven algorithms for target localization accuracy versus existing approaches, benchmarking complex-valued learning algorithms, and transfer learning in realistic radar scenarios (which enables the adaptation of pre-trained models to new environments — we demonstrate this via a toy example). Across all applications, we consider five example scenarios in RASPNet: Bonneville Salt Flats, UT ($q_{i}=1,i=29$) as the baseline scenario, Central Kansas, KS ($q_{i}=4,i=60$) from the Basic scenarios, the Ouachita Mountains, OK ($q_{i}=35,i=62$) from the Intermediate scenarios, the Great Smoky Mountains, TN ($q_{i}=57,i=76$) from the Challenging scenarios, and the Grand Canyon, AZ ($q_{i}=92,i=35$) from the Eye-watering scenarios, wherein $\smash{\tilde{S}=\{29,35,60,62,76\}}$. We first preprocess the RASPNet data via conventional RASP techniques, per Section 4.1. We compiled all results on a NVIDIA GeForce RTX 3090 GPU.
@@ 4.1 Data Preprocessing
To prepare RASPNet for the target localization and transfer learning tasks, we consider sub-regions from the scenarios, $\smash{i\in\tilde{S}}$, which we call radar processing regions. The minimum and maximum range of each radar processing region from the airborne radar platform and the azimuth angle bounds of each radar processing region are provided in Table 3 (see Appendix Section C.1), and the range bin indices are $g\in\{75,76,\ldots,95\}\subset\{1,2,\ldots,G\}$, whereby $G^{\prime}=21$. We run a set of $N^{\prime}=1000$ independent experiments, where in each experiment, a point target is placed in range and azimuth, uniformly at random inside each radar processing region. We select the radar cross section (RCS) of each target independently at random from the uniform distribution, $\mathcal{U}[15,25]$ (dBsm), and we generate $K^{\prime}=100$ independent random realizations of the radar return for each target placement. We consider stationary targets in this example, with $\Lambda^{\prime}=1$, wherein there is no Doppler processing. For compatibility with our generated target data, we divide the K=10, [^0] K=10{,}000 clutter realizations provided by RASPNet for scenario $\smash{i\in\tilde{S}}$ into $N^{*}=100$ separate clutter datasets with $K^{\prime}=100$ realizations, where $K=N^{*}K^{\prime}$. As $\Lambda^{\prime}=1$, we consider the first pulse from the clutter realizations. For the CVNN target localization benchmark task, we take the first $5$ of the $100$ realizations, where $K^{\prime}=5$. Per this construction, we now consider the signal model outlined in Section C.2 of the Appendix.
@@
For both non-CVNN target localization and transfer learning tasks, we leverage datasets of heatmap matrices (Venkatasubramanian et al., 2024), which comprise test statistics obtained via the signal model from Appendix Section C.2, where $\phi=0$. To produce each heatmap matrix, we sweep the radar steering vector, $\mathbf{\tilde{a}}(\theta,0,0)$, across azimuth with step size $\Delta\theta=0.4^{\circ}$, from $\theta_{\text{min}}$ to $\theta_{\text{max}}$, which yields an NAMF azimuth spectrum at each range bin. Arranging these spectra across the $G^{\prime}$ range bins produces a size $G^{\prime}\times(\lfloor(\theta_{\text{max}}-\theta_{\text{min}})/\Delta\theta\rfloor+1)$ matrix of NAMF test statistics. The values for $\theta_{\text{min}}$ and $\theta_{\text{max}}$ are in Table 3 for scenarios $\smash{i\in\tilde{S}}$. The cumulative dataset comprises the first 25,00025{,}000 of the $N=N^{\prime}N^{*}=100{,}000$ total heatmap matrices, formed using all combinations of $\mathbf{X}^{i}[g]$ and $(\mathbf{Z}^{i}[g]+\mathbf{A}^{i}[g])$. For the CVNN task, we consider $3$-dimensional feature vectors, which are formed from the radar return data matrices described in Section C.2 of the Appendix, where $\phi=0$. The cumulative dataset comprises the first 25,00025{,}000 feature vectors of the 100,000100{,}000 total combinations.
4.2 Target Localization Benchmark
Using the heatmap matrix datasets for scenarios $\smash{i\in\tilde{S}}$, we benchmark the target localization accuracy of two existing algorithms — the classical peak-cell midpoint algorithm outlined in (Venkatasubramanian et al., 2022) and the regression CNN architecture introduced in (Venkatasubramanian et al., 2024) — versus a new specialized architecture we call Adaptive Radar Transformer (ART). The ART applies lightweight convolutional filters to capture the local leakage pattern around the true target cell, then pools and projects these features into a compact token sequence. A transformer encoder then attends across all range–azimuth bins to resolve non‑local correlations in the leakage signature before a MLP head regresses the precise target coordinates. By adaptively weighting attention scores over leakage structures, ART can more precisely pinpoint the true target location than a purely local model can. We clarify that other classical algorithms exist for target localization, including gradient-based local search methods; these algorithms can also be benchmarked using RASPNet. Our considered model architectures are further described in Section D of the Appendix.
We now produce the training and test datasets for each scenario, $\smash{i\in\tilde{S}}$. We partition each dataset of 25,00025{,}000 heatmap matrices such that the first $N_{train}=0.8N$ examples form the training dataset, and the remaining $N_{test}=0.2N$ examples form the test dataset. Denote $(\mathbf{\bar{\text{$r$}}}\vphantom{r}_{t}^{i},\mathbf{\bar{\text{$\theta$}}}\vphantom{\theta}_{t}^{i})$ as the range and azimuth of the midpoint of the peak heatmap matrix cell, and denote $\smash{(\mathbf{\tilde{\text{$r$}}}\vphantom{r}_{t}^{i},\mathbf{\tilde{\text{$\theta$}}}\vphantom{\theta}_{t}^{i})}$ as the range and azimuth values predicted by each neural network, for example $t$ from our dataset for scenario $i$. While our heatmap matrices follow the polar coordinate system, we can transform their ground truth target locations into Cartesian coordinates to report the localization error in meters, whereby $\smash{(\mathbf{\bar{\text{$r$}}}\vphantom{r}_{t}^{i},\mathbf{\bar{\text{$\theta$}}}\vphantom{\theta}_{t}^{i})\rightarrow(\mathbf{\bar{\text{$x$}}}\vphantom{x}_{t}^{i},\mathbf{\bar{\text{$y$}}}\vphantom{y}_{t}^{i})}$ and $\smash{(\mathbf{\tilde{\text{$r$}}}\vphantom{r}_{t}^{i},\mathbf{\tilde{\text{$\theta$}}}\vphantom{\theta}_{t}^{i})\rightarrow(\mathbf{\tilde{\text{$x$}}}\vphantom{x}_{t}^{i},\mathbf{\tilde{\text{$y$}}}\vphantom{y}_{t}^{i})}$. Let $(r_{t}^{*i},\theta_{t}^{*i})\rightarrow(x_{t}^{*i},y_{t}^{*i})$ be the ground truth target location for example $t$ from our dataset. For vectors $\mathbf{s}_{t}^{*i},\mathbf{\hat{\text{$\mathbf{s}$}}}\vphantom{s}_{t}^{i}$, the localization error, $Err(\mathbf{s}_{t}^{*i},\mathbf{\hat{\text{$\mathbf{s}$}}}\vphantom{s}_{t}^{i})$, over the $N_{test}$ test examples, is:
$Err(\mathbf{s}_{t}^{*i},\mathbf{\hat{\text{$\mathbf{s}$}}}\vphantom{s}_{t}^{i})=\Big(\sum\nolimits_{t=1}^{N_{test}}\|\mathbf{s}_{t}^{*i}-\mathbf{\hat{\text{$\mathbf{s}$}}}\vphantom{s}_{t}^{i}\|_{2}\Big)/N_{test}.$ (4)
In our benchmark, $(\mathbf{s}\vphantom{s}_{t}^{*i},\mathbf{\hat{\text{$\mathbf{s}$}}}\vphantom{s}_{t}^{i})=([x_{t}^{*i},y_{t}^{*i}],[\mathbf{\hat{\text{$x$}}}_{t},\mathbf{\hat{\text{$y$}}}_{t}])$, with $(\mathbf{\hat{\text{$x$}}}\vphantom{x}_{t}^{i},\mathbf{\hat{\text{$y$}}}\vphantom{y}_{t}^{i})=(\mathbf{\bar{\text{$x$}}}\vphantom{x}_{t}^{i},\mathbf{\bar{\text{$y$}}}\vphantom{y}_{t}^{i})$ for the peak cell midpoint $Err_{\text{NAMF}}]$, and $(\mathbf{\hat{\text{$x$}}}\vphantom{x}_{t}^{i},\mathbf{\hat{\text{$y$}}}\vphantom{y}_{t}^{i})=(\mathbf{\tilde{\text{$x$}}}\vphantom{x}_{t}^{i},\mathbf{\tilde{\text{$y$}}}\vphantom{y}_{t}^{i})$ for each neural network $Err_{\text{CNN/ART}}]$. For each scenario, $\smash{i\in\tilde{S}}$, we train each neural network until convergence on the training dataset using MSE as our loss function, and we leverage the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $\alpha=1\text{e-}3$. The localization errors on the test dataset are given for the peak cell midpoint and each neural network in Table 1. We see that as the scenarios become more challenging, measured using the $\mathcal{E}$-statistic with respect to the baseline scenario, the localization performance of all methods worsen, matching our intuition from Section 3. Our proposed ART achieves SOTA performance across all scenarios.
Table 1: Benchmarking target localization performance (average Euclidean distance) for $\smash{i\in\tilde{S}}$.
[…]
1 29 21.62 $\pm$ 0 9.88 $\pm$ 0.50 7.75 $\pm$ 0.36 0
4 60 21.83 $\pm$ 0 14.60 $\pm$ 0.82 9.76 $\pm$ 0.44 3.51e-8
35 62 187.31 $\pm$ 0 68.31 $\pm$ 2.23 51.86 $\pm$ 1.59 2.63e-7
57 76 360.53 $\pm$ 0 145.34 $\pm$ 5.91 123.08 $\pm$ 2.40 7.91e-7
92 35 1004.20 $\pm$ 0 220.45 $\pm$ 9.67 180.14 $\pm$ 5.50 4.04e-6
@@ 4.2 Target Localization Benchmark
Figure 4: Test performance comparison on RASPNet scenarios $i\in\tilde{S}$ using convolutional CVNN, convolutional RVNN, and convolutional Steinmetz and analytic networks [ours]. The x-axis represents the training epochs, while the y-axis indicates the target localization performance.
4.3 CVNN Target Localization Benchmark
We now benchmark the performance of CVNNs, recalling the datasets of feature vectors described in Section 4.1 for scenarios $\smash{i\in\tilde{S}}$. Extending (Trabelsi et al., 2017; Venkatasubramanian et al., 2025), we compare the performance of convolutional CVNNs and real-valued neural networks (RVNNs), versus Steinmetz and analytic neural networks (shown to be SOTA for complex-valued learning tasks) via a new Steinmetz variant called Convolutional Steinmetz Networks for improved spatial processing of complex data. We use the partitioning scheme from Section 4.2 and compute localization errors via Eq. (4). For each scenario, $\smash{i\in\tilde{S}}$, we train all networks until convergence on the training dataset using MSE Loss and the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $\alpha=5\text{e-}2$. Per Figure 4, while the convolutional RVNN and CVNN falter, the convolutional Steinmetz and analytic networks effectively learn the task across all scenarios $\smash{i\in\tilde{S}\setminus\{35\}}$, achieving SOTA performance.
@@ 4.4 Transfer Learning
Apart from benchmarking the performance of algorithms, RASPNet can also be leveraged for transfer learning applications in realistic adaptive radar processing scenarios. We recall that the RFView®-generated data exhibits some degree of statistical similarity to real-world data, as examined in Section 2.4. Now, suppose we collect $K$ clutter realizations, $\{z_{j}^{\omega},\ \forall j\in D\}$ from some real-world scenario, $\omega$, where $z_{j}^{\omega}\sim Z^{\omega}$, and train $100$ models for some predictive adaptive radar processing task on each of the $100$ RASPNet scenarios, $i\in S$. Our aim is to find the trained model that has the best performance for the same task on the real-world scenario, $\omega$. We can proceed by ranking the RASPNet scenarios, $i\in S$, by their similarity to scenario $b=\omega$, using Algorithm 2. We hypothesize that the scenario with the lowest $\mathcal{E}$-statistic, relative to the real-world scenario, is the optimal choice for selecting the corresponding trained model. To improve the performance of this trained model on the real-world scenario, we can use a few samples from the real-world scenario to fine-tune it for the task at hand.
Figure 5: Regression CNN target localization performance for scenarios $\smash{i\in\tilde{S}\setminus\{29\}}$ after transfer learning $([Err_{\text{CNN}}]_{\text{TL}})$ and fine-tuning $([Err_{\text{CNN}}]_{\text{FT}})$.
As a proof-of-concept, we consider again the target localization task and train the regression CNN to localize targets in scenario $i=29$ (Bonneville Salt Flats, UT). However, we evaluate this trained network on each scenario $\smash{i\in\tilde{S}\setminus\{29\}}$ to evaluate whether the scenario with the highest similarity to $i=29$ from $\smash{\tilde{S}\setminus\{29\}}$ achieves the best localization performance. As in Section 4.2, our training dataset comprises the first $0.8N$ of the 25,00025{,}000 heatmap matrices generated for scenario $i=29$. We consider four test datasets, each comprising the last $0.2N$ of the 25,00025{,}000 total heatmap matrices from scenario $\smash{i\in\tilde{S}\setminus\{29\}}$, and we consider four datasets for fine-tuning, each comprising the first $64$ of the 25,00025{,}000 total heatmap matrices from scenario $\smash{i\in\tilde{S}\setminus\{29\}}$. We train the regression CNN on the training dataset using MSE Loss, and we leverage the Adam optimizer (Kingma and Ba, 2014) with $\alpha=1\text{e-}3$. The localization accuracies on the test datasets, $Err_{\text{CNN}}]_{\text{TL}}$, $\smash{\forall i\in\tilde{S}\setminus\{29\}}$, are depicted in Figure 5. As expected, the best performance is given by the model evaluated on the scenario $i=60$ (Central Kansas). This scenario has the lowest $\mathcal{E}$-statistic with respect to $i=29$ from $\smash{\tilde{S}\setminus\{29\}}$.
We now consider the case of fine-tuning the trained regression CNN using a small set of $64$ heatmap matrices from the scenarios $\smash{i\in\tilde{S}\setminus\{29\}}$. We freeze the convolutional and batch normalization layers of our trained regression CNN, halving the number of parameters. Using the $64$ heatmap matrices, the weights and biases of the remaining layers are fine-tuned using MSE Loss and the Adam optimizer with $\alpha=1\text{e-}3$. After fine-tuning the regression CNN for each scenario $\smash{i\in\tilde{S}\setminus\{29\}}$, we obtain the localization accuracies on the test datasets, $Err_{\text{CNN}}]_{\text{FT}}$, $\smash{\forall i\in\tilde{S}\setminus\{29\}}$, depicted in Figure 5. We note that the improvements afforded by the regression CNN are largely recovered after fine-tuning.